A proton moving to the right at 6.2 × 105 ms-1 enters a region where there is an electric field of 62 kNC-1 directed to the left. Describe qualitatively the motion of the proton in this filed. What is the time taken by the proton to come back to the point where it entered the field? (Use the standard values of the mass and charge of a proton)
Answer #1

acceleration acts on proton

a = qE/ m

a = 1.6* 10^-19* 62* 1000 / ( 1.67* 10^-27)

a = 5.94* 10^12 m/s^2

now time it takes to again come to the same point of insertion

t = 2 v/a

t = 2 * 6.2* 10^5 / (5.94* 10^12)

t = 2.0875* 10^-7 s

=======

Comment before rate in case any doubt, will reply for sure.. goodluck